OpenStudy (s.gracex):

What is the solution to the system of equations? Use the substitution method. y+2x=7 14-4x=2y A. the solution is (1,5) B. the solution is (21,0) C. there is an infinite number of solutions D. There is no solution

OpenStudy (anonymous):

Try and use substitution. Here is an example: The substitution method is a method that is based off a rule of algebra. Basically, it's similar to using a formula. You substitute values into the formula. The same thing applies here. Let's say we have the following system: y = 2x + 3 5x + 2y = 24 Before we turn over to elimination, let's try using our new friend substitution. It clearly says that y is equal to 2x + 3. This means that we can substitute 2x + 3 in for y. 5x + 2y = 24 5x + 2(2x + 3) = 24 5x + 4x + 6 = 24 Now that we have eliminated a variable, we can solve the system. 9x + 6 = 24 9x = 18 x = 2 Now let's plug this back into one of the equations. I prefer to plug it into the equation that has a variable isolated. y = 2x + 3 y = 2(2) + 3 y = 4 + 3 y = 7 Therefore, the solution set is (x, y) = (2, 7).

OpenStudy (anonymous):

Now try and do it yourself, you can follow the guidelines :)